OpenStudy (anonymous):

P(sum of 6 and both numbers even)

OpenStudy (anonymous):

if you roll a pair of dice?

OpenStudy (anonymous):

two fair dice

OpenStudy (anonymous):

then we can list the two ways to get a 6 with both even \((2,4)\) and \((4,2)\)

OpenStudy (anonymous):

my answer was 1/9 but i wasnt sure if i was right

OpenStudy (anonymous):

there are \(6\times 6=36\) possible rolls two are favourable, so no, it not \(\frac{1}{9}\) it is \(\frac{2}{36}=\frac{1}{18}\)

OpenStudy (anonymous):

thank you there are two more parts to the question. what are the odds of getting a smaller number on the second roll. P( same number or sum greater than 8)